# Any-Order Flexible Length Masked Diffusion

### v1

@@ F.3 Weight Initialization training from LLaDA
In this section, we describe the procedure for adapting the pretrained LLaDA-8B base model into the FlexMDM, a natural choice given that the design of FlexMDM builds incrementally upon the design of MDM.
@@
LLaDA uses a bidirectional transformer, relying on the fact that the MDM modeling does not require an explicit time embedding [Zheng et al., 2024]. For FlexMDM, we inject a time-embedding layer via AdaLN [Peebles and Xie, 2023], which is necessary to model insertion expectation. For parameter efficiency, we tie (share) the four AdaLN parameter sets across the intermediate Transformer layers. We also attach a softplus scalar head to model the insertion expectation.
We then attach the LoRA adapter [Hu et al., 2022] at every attention(q_proj, k_proj, v_proj) as well as the unmasking posterior head. This is because MDM and FlexMDM’s task to learn the unmasking posteriors are different, as it is used for a semantically different task in FlexMDM, where the unmasking has to account for the shifting of tokens due to insertion. We highlight that this choice of fine-tuning the last head is crucial to obtain readable variable-length generations.
@@
Throughout, LoRA uses $r=128$, $\alpha=128$, and dropout $0.1$. Training runs for 200,000200{,}000 gradient steps with batch size $64$ on $16$ H100s over three days. For optimization, we use AdamW [Loshchilov and Hutter, 2017] with learning rate $1\times 10^{-4}$ and weight decay $0.1$, together with a cosine warm-restarts scheduler.
Instruction-fine tuning.
For infilling, both models are instruction finetuned for 50 epochs on the educational-instruct split of opc-sft-stage2. In training, the model is given the prefix and the suffix, and is asked to predict insertion between the prefix and suffix. Formally, this corresponds to using a data-dependent coupling Albergo et al. [2024].
For GSM8k, FlexMDM is instruction finetuned for 1000 epochs while LLaDA is finetuned for 100 epochs. The difference in the number of iterations is due to the observation that the performance of LLaDa plateaus under a smaller number of training iterations while FlexMDM continues to improve. We conjecture this is due to the fact the pretraining dataset LLaDa use contain a large amount of mathematical reasoning text, while the unmasking head for FlexMDM might have been undertrained on mathematical tasks during the adaption phase from LLaDa to FlexMDM.
In both cases, hyperparemeters are chosen to be identical as pretraining.
Inference.
For FlexMDM, we adopt the Top-K probability confidence-based sampling with sliding window with $\gamma_{1}=5.0,\gamma_{2}=64$, for details of the algorithm see Appendix E. For MDM, we report the best of a semiautoregressive confidence-based sampling strategy and confidence-based strategy following Nie et al. [2024]. In the case for FlexMDM, We highlight that using a sliding window significantly improves performance, which is surprisingly not the case for MDM with semi-autoregressive sampling. In general, using confidence-based sampling also greatly improves over vanilla inference. For both math and coding, low temperature sampling is crucial in obtaining strong Pass@1 performance, we use temperature equal to $0.0$, equivalently this can be seen as always unmasking towards the most likely token. Similar trends have been observed in prior work Nie et al. [2024]. If one is to obtain competitive Pass@k performance $k>0$, prior work Cobbe et al. [2021] has observed that using higher temperature can be beneficial, we did not explore such choices in this paper

### v2

@@ F.3 Weight Initialization training from LLaDA
In this section, we describe the procedure for adapting the pretrained LLaDA-8B base model into the FlexMDM.
@@
LLaDA uses a bidirectional Transformer without an additional time embedding, leveraging the fact that MDM does not require an explicit time signal [Zheng et al., 2024]. For FlexMDM, to model the insertion expectation, we inject a time-embedding pathway via AdaLN [Peebles and Xie, 2023]. For parameter efficiency, we tie (share) the four AdaLN parameter sets across the intermediate Transformer layers. We also attach a softplus scalar head to parameterize the insertion expectation.
Next, we attach LoRA adapters [Hu et al., 2022] to every attention projection (q_proj, k_proj, v_proj) and to the unmasking-posterior head. We include LoRA on the unmasking posterior head because the unmasking posteriors differ between MDM and FlexMDM: in FlexMDM, unmasking must account for token shifts induced by insertions. This fine-tuning of the last head is crucial for enabling variable-length generation.
@@
The LoRA configuration that we use is $r=128$, $\alpha=128$, and dropout $0.1$. Training runs for $200,000$ gradient steps with a batch size of $64$ on $16$ H100s, which took approximately 3 days. We optimize with AdamW [Loshchilov and Hutter, 2017] at learning rate $1\times 10^{-4}$ and weight decay $0.1$, using a cosine warm-restarts scheduler.
Evaluation on GSM8K.
We instruction-fine-tune (IFT) the FlexMDM base model on the GSM8K training split. To preserve the instruction–answer format, we modify the interpolant in Eq. 6 so that tokens corresponding to the instruction are excluded from the interpolant—these tokens remain fixed for all time steps. We apply the same strategy to obtain the baseline (that is, IFT from LLaDA-base), modifying the MDM interpolant so that instruction tokens remain fixed. This choice is equivalent to the IFT recipe used in Nie et al. [2025], Ye et al. [2025]. Both models are IFT-ed for $1000$ epochs. (Other IFT hyperparameters match those used in our base setup unless otherwise noted.)
For FlexMDM inference, we start from the instruction prompt at $t=0$ and run adaptive inference to $t=1$. Concretely, we use Top-K probability with a sliding window (Appendix E) with $\gamma_{1}=5.0$ and $\gamma_{2}=64$. For LLaDA, we report the best result under the semi-autoregressive, confidence-based sampling of [Nie et al., 2025]. For both models, we set the token sampling temperature to $0.0$, which we confirm to be important for strong Pass@1. Overall, adaptive inference substantially improves performance over vanilla inference.
Evaluation on HumanEval-infill.
Code infilling conditions on a prefix and suffix, and asks the model to complete the middle part of the code. For FlexMDM, we format training examples as $(\texttt{Prefix};\texttt{[SEP]};\texttt{[Middle]};\texttt{[SEP]};\texttt{Suffix})$, where [SEP] is a separator token and Instruction describes the infilling task for a model. We modify Eq. Eq. (6) so that tokens for Prefix, Suffix, and [SEP] are fixed in the interpolant (i.e., excluded from the interpolant). Thus, at $t=0$ the state is $(\texttt{Prefix};\texttt{[SEP]};\texttt{[SEP]};\texttt{Suffix})$.
For MDM, we use the format $(\texttt{Instruction};\texttt{[PRE]};\texttt{Prefix};\texttt{[SUF ]};\texttt{Suffix};\texttt{[SEP]};\texttt{Middle})$, with Instruction prompting infill after prefix and suffix , along with [PRE] and [SUF] separate the prefix and suffix, respectively. Here too, the tokens without Middle are held fixed by the modified interpolant. This difference in formatting reflects the fixed-length nature of MDMf MDM (no token insertion). This formatting has been used in the code infilling tasks for autoregressive models in Bavarian et al. [2022]. Naively masking the Middle span yields a base state at $t=0$ of $(\texttt{Prefix};\texttt{Masked};\texttt{Suffix})$, where Masked is a fully masked sequence of length $|\texttt{Middle}|$. This leaks length information—materially simplifying the task— and renders comparisons to FlexMDM unfair, since FlexMDM does not observe the target span length. For fair evaluation, we therefore avoid length-revealing masks and require methods to infer the span length during inference.
We IFT both models on the educational-instruct split of opc-sft-stage2; the architecture and optimization configurations match those used for GSM8K IFT. At evaluation, we initialize from the base distributions: for FlexMDM, $(\texttt{Instruction};\texttt{Prefix};\texttt{[SEP]};\texttt{[SEP]};\texttt{Suffix})$; for MDM, $(\texttt{Instruction};\texttt{Prefix};\texttt{Suffix};\texttt{[SEP]})$. We use the same Top-K adaptive inference for both and temperature $0.0$. Final outputs are scored with the HumanEval-infill verifier toolkit to compute success rates.